OpenStudy (anonymous):

Does the following function have any factors at all? f(x) = -2x^4 + 7x^2 - 3

OpenStudy (saifoo.khan):

We suppose x^2 = y So we will get, -2y^2 + 7y - 3 Can you factor now?

OpenStudy (anonymous):

\[-2x^4+7x^2-3=-2x^4+6x^2+x^2-3\]\[=-2x^2(x^2-3)+(x^2-3)=(x^2-3)(-2x^2+1)\]

OpenStudy (saifoo.khan):

@joemath314159 , is my method correct as well?

OpenStudy (anonymous):

yes. Your method makes it a little easier to see what to do. People are generally comfortable with quadratics.

OpenStudy (lgbasallote):

i like joe's method more :P haha jk

OpenStudy (saifoo.khan):

HAHA @lgbasallote

OpenStudy (saifoo.khan):

Btw @joemath314159 , how are you? It's been long since you been here! :D

OpenStudy (anonymous):

Im doing good :) only one more school week left in the semester. Things are winding down.

OpenStudy (saifoo.khan):

Cool. xD

OpenStudy (anonymous):

Joe how did you get to that method?

OpenStudy (anonymous):

it was something I learned here a while back. Its called the "a-c" method i think. when trying to factor something like:\[ax^2+bx+c\]you multiply a and c, and think of two numbers that multiply to ac, but add up to b. In our case, we needed to think of two numbers that multiply to 6, but add up to 7, which is 6 and 1. Notice how I split the 7x^2 into 6x^2+1x^2.

OpenStudy (anonymous):

Got it. Thanks

OpenStudy (anonymous):

But, there is one thing I do not understand. When you have a composite function, is it correct to say that it its domain is determined by: range(outer fcn) intersect domain(inner fcn)?

OpenStudy (anonymous):

i think it might be the other way around. The range of the inner function intersect the domain of the outer...i think.

OpenStudy (anonymous):

wait wait..but you have to watch the domain of the inner as well. hmm..

OpenStudy (anonymous):